### v1

# From a Tiny Slip to a Giant Leap: An LLM-Based Simulation for Fake News Evolution

Abstract.
With the growing spread of misinformation online, research has increasingly focused on detecting and tracking fake news. However, an overlooked issue is that fake news doesn’t naturally exist in social networks—it often originates from distorted facts or deliberate fabrication by malicious actors. Understanding how true news gradually evolves into fake news is critical for early detection and prevention, reducing its spread and impact. Hence, in this paper, we take the first step toward simulating and revealing this evolution, proposing a Fake News evolUtion Simulation framEwork (FUSE) based on large language models (LLMs). Specifically, we employ LLM as agents to represent individuals in a simulated social network. We define four types of agents commonly observed in daily interactions: spreaders, who propagate information; commentators, who provide opinions and interpretations; verifiers, who check the accuracy of information; and bystanders, who passively observe without engaging. For simulated environment, we model various social network structures, such as high-clustering networks and scale-free networks, to mirror real-world network dynamics. Each day, the agents engage in belief exchanges, reflect on their thought processes, and reintroduce the news accordingly. Given the lack of prior work in this area, we developed a FUSE-EVAL evaluation framework to measure the deviation from true news during the fake news evolution process. The results show that FUSE successfully captures the underlying patterns of how true news transforms into fake news and accurately reproduces previously discovered instances of fake news, aligning closely with human evaluations. Moreover, our work provides insights into the fact that combating fake news should not be delayed until it has fully evolved; instead, prevention in advance is key to achieving better outcomes. We hope our work inspires further research and responsibility in the early detection and prevention of fake news in social networks.  [^1] Our code and data will be released in camera ready version.
Fake News Evolution, LLM-Based Simulation, Agent, Intervention
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems World Wide Webccs: Computing methodologies Artificial intelligence
1. INTRODUCTION
The rapid spread of fake news has become a significant global concern, with social media accelerating the production, consumption, and propagation (Lazer et al., 2018b; Olan et al., 2022). Addressing the evolution of fake news is therefore critical to mitigating its far-reaching effects and is particularly important for fostering a responsible web environment (Aïmeur et al., 2023; Qin, 2020; Adriani, 2019). Most existing studies on fake news focus on detecting misinformation or simulating its spread once it has already been generated (Garimella et al., 2017; Wang et al., 2019). For instance, Piqueira et al. (2020) categorized individuals into four types and used mathematical models to simulate the spread of fake news, as depicted in Figure 1(a), showing the dynamics of different actor populations over time. On a micro-level, Jalili and Perc (2017) defined numerical conditions for opinion change to study fake news dissemination, as shown in Figure 1(b). The above work follows a key assumption common in much of the existing research: fake news inherently exists within social networks.
However, a significantly overlooked issue in these works is that fake news does not naturally exist within social networks. In reality, it may originate from a piece of true news that, over time, becomes distorted, misinterpreted, or altered as it spreads among individuals, eventually evolving into fake news (Guo et al., 2021; Wang et al., 2021; Sunstein, 2014; Herzing and Muck, 2024; Shen et al., 2024) as illustrated in Figure 1(c), or it may be intentionally generated by malicious actors, to grab public attention or serve specific agendas such as political manipulation or financial gain (Kimmel, 2013; Schindler, 2007). Here we define partially evolved fake news as news content that contains elements of factual information but has been manipulated, misrepresented, or selectively edited, resulting in a mixture of both accurate and false details. The dynamic and evolving nature of fake news underscores the need for a deeper understanding of the underlying transformation mechanisms. Addressing this gap is crucial, as understanding these mechanisms can help us design more effective early interventions to prevent fake news from gaining traction within social networks (Schindler, 2007; Shen et al., 2024; DiFonzo and Bordia, 2011).
Figure 1. (a) Macro-level observation of population dynamics based on the mathematical model, categorizing individuals into four types and showing their quantity changes over time. (b) Micro-level visualization of the conventional fake news dissemination model, which assumes fake news inherently exists. (c) Micro-level evolution of fake news, where true news gradually transforms into fake news as it propagates through a network, with content alterations at various stages.
Correspondingly, in this paper, we focus on modeling and understanding fake news formative process, and present the first Fake news evolUtion Simulation framEwork (FUSE). Unlike traditional epidemiological models, which assume predefined transition probabilities between different states of belief (Nekovee et al., 2007; N Zehmakan and Galam, 2020), we employ LLMs to simulate human behaviors, generate sophisticated interactions, reflect on information, and give their own version of news report. For the simulation environment setting, we construct various social network structures, such as high-clustering and scale-free networks, to mirror real-world online interaction patterns. Under this framework, we define four types of agents commonly observed in daily interactions: spreaders, who propagate information; commentators, who provide opinions and interpretations; verifiers, who check the accuracy of information; and bystanders, who passively observe without engaging. Each agent represents an individual with unique attributes, including name, age, personality traits, and education level. Each day, agents interact with its neighbor peers, exchanging beliefs. They adjust their beliefs based on their reasoning, prior experiences, and the views of others. Agents are equipped with hierarchical memory: short-term memory records daily interactions, while long-term memory stores accumulated knowledge and prior beliefs. Additionally, agents use a reflective reasoning process, enabling them to reconsider past experiences and update their believed news content dynamically.
To our best knowledge, there is no prior work on language-based evaluation of fake news evaluation. Hence, we introduce FUSE-EVAL, which tracks information distortion and provides quantitative insights into how true news evolves into fake news. It is a multi-dimensional evaluation framework that analyzes the degree of deviation between news after evolution and the original news, which includes aspects such as Sentiment Shift (SS), New Information Introduced (NII), Certainty Shift (CS), STylistic Shift (STS), Temporal Shift(TS), and Perspective Deviation (PD). Our FUSE framework is also validated through comprehensive simulation experiments, demonstrating a strong alignment with real-world observations from prior research. First, our simulations reveal a clear accumulation distortion effect as news spreads through the network, with information gradually deviating further from its original form (De Fleur, 1962; de Paula et al., 2024; Nian et al., 2022). Additionally, in high-clustering networks, true news evolves into fake news more rapidly compared to scale-free and random networks, consistent with previous studies (Trpevski et al., 2010; Lind et al., 2007). Furthermore, political fake news spreads significantly faster than topics such as terrorism, natural disasters, science, urban legends, or financial information, which also aligns with earlier findings (Rubin, 2022; Lazer et al., 2018a).
To construct a responsible web environment, our research underscores the importance of strategic interventions during the early stages of fake news evolution. Rather than waiting until misinformation has widely spread, we introduce an official agent that actively intervenes when the degree of information deviation reaches a critical threshold. This agent issues timely official statements, incorporating reliable sources to slow down or counteract the evolution of fake news. By providing a trusted narrative early in the process, this approach helps correct circulating misinformation, reducing its impact and limiting further spread. These findings offer practical insights into the role of timely, authoritative interventions for effective misinformation governance.
Our contribution can be summarized in the following ways:
  • We propose FUSE, an LLM-based simulation framework to investigate how true news gradually evolves into fake news, considering the roles of different types of agents and various social network structures observed in real-world.
  • We introduce FUSE-EVAL, a novel evaluation framework to measure the deviation from true information during the news evolution process.
  • We conduct experiments to show that our simulation aligns with real-world observations and previous studies.
  • We propose and evaluate multiple intervention strategies aimed at mitigating the spread of fake news during its evolution.
2. RELATED WORK
2.1. Fake News Evolution
Recent research into fake news evolution has focused on how misinformation spreads and transforms over time. Zhang et al. (2013) found that rumors evolve as they are repeatedly modified, becoming shorter and more shareable, while Guo et al. (2021) empirically tracked fake news evolution, noting how sentiment and text similarity change as truth transitions into misinformation. Xia et al. (2020) developed a sentiment analysis pipeline that addresses sarcasm and out-of-domain input, helping track shifts in public opinion related to fake news. Other studies have emphasized structural and behavioral aspects of fake news propagation. Zhao et al. (2024) proposed a dynamic learning method that captures temporal changes in rumor propagation, revealing how rumor patterns evolve. Wang et al. (2021) demonstrated that slight content changes during the COVID-19 pandemic aided the spread of misinformation, while Li et al. (2016) examined how user behaviors, particularly the role of verified accounts, influence the evolution and dissemination of rumors. These studies underscore the dynamic nature of fake news evolution and the need for comprehensive approaches.
However, there has not been a detailed and comprehensive study on how true news evolves into fake news, with only some superficial linguistic analyses (Zhang et al., 2013; Guo et al., 2021).
2.2. Simulations of Fake News Propagation
Fake news propagation modeling plays a vital role in understanding the dynamics of misinformation spread and is key to developing intervention strategies such as early warning systems (Garimella et al., 2017), misinformation blocking (Song et al., 2015), and truth confrontation (Wang et al., 2019). Current models can be broadly categorized into three types: epidemic models, point process-based models, and diffusion models (Sun et al., 2023). Epidemic models, inspired by the spread of infectious diseases, categorize users as susceptible, infected, or recovered. The SIR model (Zhu and Wang, 2017) is widely used to simulate how misinformation spreads and shows how network structures affect its speed and reach (López et al., 2024). Point process-based models treat fake news spread as random events over time, capturing the temporal dynamics and predicting key influencer engagement (Chen et al., 2019; Gao et al., 2019). Diffusion models focus on individual thresholds, where users decide to spread misinformation based on peer influence, making them effective for studying echo chambers and tight-knit communities (Jalili and Perc, 2017).
Our work extends these models, using LLMs to simulate the evolution of true news into fake news.
2.3. LLM-based Agents for Social Simulation
Agent-based modeling and simulation focuses on representing complex systems by simulating individual agents and their interactions within a dynamic environment (Macal and North, 2005). The integration of LLMs into agent-based modeling has significantly expanded the potential of such simulations, particularly in digital environments where LLM agents excel in natural language tasks (Chen et al., 2023b, a; Liu et al., 2024a). Moreover, LLM-based agents can simulate human-like intelligence, actively engaging in tasks through self-directed planning and decision-making (Xi et al., 2023), and their capabilities extend beyond text generation (Yuhan et al., 2023). These strengths have led to the broad adoption of LLM agents across various domains (Liu et al., 2024b; Park et al., 2022; Zhang et al., 2024; Li et al., 2023; Park et al., 2023; Lin et al., 2023; Kovač et al., 2023; Jinxin et al., 2023; M. Bran et al., 2024; Boiko et al., 2023), making them a new paradigm for simulation by embedding human-level intelligence into agents.
In more specific applications, LLM agents have been employed to simulate social media dynamics. For instance, Törnberg et al. (2023) used LLMs with agent-based modeling to investigate social media algorithms and provide insights into real-world phenomena. Similarly, Park et al. (2022) demonstrated that LLM agents could generate social media content indistinguishable from human-generated material. These developments showcase the growing potential of LLM agents in modeling human social behavior at a collective level. To the best of our knowledge, our work is one of the first to apply LLMs in simulating the evolution of fake news within an agent-based framework.
3. METHODOLOGY
3.1. Problem Formulation
In this study, we aim to simulate the gradual evolution of true news into fake news within a social network by leveraging LLMs as agents. We construct a simulation environment with a pool of $N$ agents $\mathcal{A}=(a_{1},\ldots,a_{N})$, each endowed with a unique persona that includes their role type (spreader, commentator, verifier, or bystander), personality traits, demographic information, and initial opinion toward a news topic.
At time $t=0$, a piece of true news $S_{0}$ is introduced into the network. The agents are connected according to a predefined social network structure, $\mathcal{G}=(\mathcal{A},\mathcal{E})$, which may represent high-clustering, scale-free, or random networks to reflect real-world dynamics. On each day $t=1,2,\ldots,T$, agents interact with their neighbors as defined by the network structure. During these interactions, they exchange information and opinions, influenced by their personas and prior knowledge. After interactions, agents reintroduce the news content considering the new information received. The evolution of the news content held by agent $a_{i}$ at time $t$, denoted as $S_{i}^{t}$, depends on their previous news content and the information received from their neighbors. This update process is defined by:
(1) $S_{i}^{t}=f(S_{i}^{t-1},\{S_{j}^{t-1}|a_{j}\in\mathcal{N}_{i}\},\mathcal{P}_{i}),$
where $f(\cdot)$ represents the agent’s information processing function, influenced by their persona $\mathcal{P}_{i}$ and the interactions with neighboring agents $\mathcal{N}_{i}$.
Figure 2. Our FUSE framework simulates news evolution by equipping each agent with role-based decision-making capabilities. Propagation Role-aware agents (PRA) process true news through interactions within the news evolution simulator (NES), where their role identities shape how they engage with the news.
Our objective is to analyze how the true news $S_{0}$ transforms over time due to the agents’ interactions and personal biases. This simulation allows us to research the roles of different agent types, network structures, news topics, and individual personality traits in evolution. By tracking the changes in $S_{i}^{t}$, we can quantify the deviation from the original news and understand the factors influencing the transformation into fake news.
3.2. Our Simulation Framework
As depicted in Figure 2, our FUSE framework integrates two core components: the Propagation Role-Aware agents (PRA) and the News Evolution Simulator (NES). The PRA module equips each agent with role-based decision-making capabilities, enabling them to process and disseminate information according to their assigned roles and personal attributes. The NES constructs the interaction environment, simulating the social network through which news propagates and evolves. Within the PRA module, each agent is powered by an LLM and is assigned a specific role type along with attributes. It influences how agents interpret information, interact with others, and update their opinions. Daily interactions are facilitated by the NES, which determines which agents interact based on the predefined social network structure, $\mathcal{G}=(\mathcal{A},\mathcal{E})$. The NES simulates various network types. During each day, agents engage with their neighbors, exchanging news content and opinions influenced by their roles and personal attributes. In scenarios where interventions are introduced—such as official announcements after the deviation level of news reaches a set threshold, aiming to correct misinformation and release credible information sources to relevant agents. At the end of each day, the simulation progresses by a one-time step, and the agents’ states are updated.
3.3. Propagation Role-Aware Agent
The PRA is the core component of our FUSE framework, designed to simulate individual human behaviors in the context of news evolution. By equipping agents with specific roles and personal attributes, we aim to closely mimic the diversity and complexity of human interactions in social networks.
3.3.1. Personal Information.
[…]
3.3.2. Role-Specific Behaviors.
At each time step $t_{i}$, agent $a_{i}$ holds a version of the news content $S_{i}^{t}$. When interacting with neighboring agents $\mathcal{N}_{i}$ as defined by the network $\mathcal{G}$, agent $a_{i}$ receives news content $\{S_{j}^{t-1}|a_{j}\in\mathcal{N}_{i}\}$. The agent then reintroduces news based on their role and persona through a role-specific update function:
(2) $f_{role}=f_{r_{i}}(S_{i}^{t-1},\{S_{j}^{t-1}|a_{j}\in\mathcal{N}_{i}\},\mathcal{P}_{i}).$
For the role in our model, spreaders may combine and amplify sensational aspects of the news, commentators may add personal interpretations or opinions, verifiers may cross-reference information before accepting or sharing it, and bystanders typically retain their previous news content $S_{i}^{t}=S_{i}^{t-1}$ unless significantly influenced (Sun et al., 2023). The prompt for the role-specific reintroduction function $f_{r_{i}}$ is in Appendix A.
3.3.3. Memory and Reflection.
In our simulation, agents engage with their neighbors each day, leading to an updated version of the news. However, owing to the potentially vast volume of interactions, storing all of them in detail is impractical. Therefore, agents are equipped with a hierarchical memory system comprising short-term memory (STM) $M_{i}^{S}$ and long-term memory (LTM) $M_{i}^{L}.$ The STM stores recent interactions and immediate information received, while the LTM accumulates knowledge over time, influencing future information processing and news updates. After interactions, agents reflect and reintroduce the news through a memory consolidation function:
(3) $M_{i}^{L,t}=g(f_{L}(M_{i}^{L,t-1}),f_{S}(M_{i}^{S,t})),$
where $g(\cdot)$ integrates new information into LTM, allowing agents to exhibit dynamic behaviors such as gradually changing their opinion on a topic or reinforcing existing opinions. The function prompt for short-term memory $f_{S}$ and long-term memory $f_{L}$ is in Appendix A.
3.3.4. Decision-Making Process.
In our FUSE framework, each agent’s opinion evolves through a reasoning process influenced by their role, persona, and interactions with others. After daily interactions and memory updates, agents reflect on the news content they hold, leading to gradual changes in their opinions. This process integrates the agent’s role in propagation—whether they are a spreader, commentator, verifier, or bystander—with their cognitive functions. The decision-making process for agent $a_{i}$ at time $t$ is modeled as:
(4) $S_{i}^{t}=f_{dm}(S_{i}^{t-1},m_{i}^{L,t-1},r_{i},\mathcal{P}_{i}).$
The function prompt $f_{dm}$ is in Appendix A, which captures how agents integrate new information with their existing opinions, taking into account their role in the decision-making process. For example, the reasoning of spreaders may lead to greater changes in $S_{i}^{t}$. Commentators often add subjective nuances to the news content through their reasoning. Verifiers are aiming to correct inaccuracies and potentially reduce misinformation. Bystanders have minimal engagement; their reasoning leads to minor or no changes in their held content.
Algorithm 1 FUSE Framework for Fake News Evolution
1:  Input: Number of agents $N$, total simulation days $T$, social network structure $\mathcal{G}=(\mathcal{A},\mathcal{E})$, original news content $S_{0}$
2:  Output: Final news content $S_{i}^{T}$ and final memory states $M_{i}^{L,T}$ for each agent $a_{i}$
3:  Initialize propagation role-aware agents:
4:  for each agent $a_{i}$ in $1$ to $N$ do
5:     Assign a propagation role $r_{i}$ and persona profile $\mathcal{P}_{i}$
6:     Set initial news content $S_{i}^{0}=S_{0}$
7:     Define short-term memory $M_{i}^{S,0}$ and long-term memory $M_{i}^{L,0}$
8:  end for
9:  Simulate daily news evolution:
10:  for each day $t$ in $1$ to $T$ do
11:     for each agent $a_{i}$ do
12:        Select neighbors $\mathcal{N}_{i}$ based on the network structure $\mathcal{G}$
13:        Receive news content $\{S_{j}^{t-1}|a_{j}\in\mathcal{N}_{i}\}$
14:        Update short-term memory $M_{i}^{S,t}$ for agent $a_{i}$ with details from the day’s interactions
15:        Based on $M_{i}^{S,t}$, update long-term memory $M_{i}^{L,t}$ for agent $a_{i}$ using Equation (3)
16:        Agent $a_{i}$ reintroduce news content $S_{i}^{t}$ using Equation (4)
17:     end for
18:  end for
19:  return  Final news content $S_{i}^{T}$ and long-term memory $M_{i}^{L,T}$ for each agent $a_{i}$
3.4. News Evolution Simulator
The News Evolution Simulator (NES) serves as the core environment where news content propagates and evolves over time through interactions among agents. This module simulates the flow of information within a social network structure $\mathcal{G}=(\mathcal{A},\mathcal{E})$, enabling the study of how true news gradually transforms into fake news due to the influence of agent behaviors, memory systems, and social interactions.
NES models a dynamic, evolving social network where agents interact based on predefined topologies. Specifically, the framework explores network structures such as random networks, where the edges $\mathcal{E}$ between agents $a_{i}\in\mathcal{A}$ are formed randomly. This simulates environments where individuals are not tightly clustered. In scale-free networks, certain agents, or hubs, have significantly more edges, simulating influential actors or ”super-spreaders” in social media. High-clustering networks are modeled to form communities where agents are highly connected within their local group, reflecting real-world interaction patterns seen in tightly-knit social circles. The network structure $\mathcal{G}$ dictates which agents interact daily, influencing the rate and manner in which news spreads and evolves in the system. The overall algorithm is shown in Algorithm 1.
3.4.1. Intervention Mechanisms
A critical feature of the NES is its ability to simulate interventions aimed at countering the evolution of fake news, thereby contributing to the development of a more responsible web environment. When the degree of deviation between the current news content $S_{i}^{t}$ and the original news $S_{0}$ surpasses a predefined threshold, an official agent is introduced into the system to correct misinformation. This official agent represents authoritative sources that provide verified information, to mitigate the spread and impact of distorted or false narratives.
The intervention process starts with continuous monitoring of the deviation between the news content held by each agent and the original news. Once the deviation exceeds a critical threshold, the official agent is triggered to take action. This agent issues official announcements based on reliable sources, targeting agents most likely to propagate or exacerbate misinformation. The prompt for the official agent is in Appendix A.
4. FUSE-EVAL: News Evolution Analysis
To systematically measure how true news evolves into fake news within our simulation, we propose a comprehensive evaluation framework named FUSE-EVAL. This framework consists of two sets of metrics: Content Deviation Metrics and Statistical Deviation Metrics, which together provide a detailed understanding of how misinformation evolves within the simulated environment.
4.1. Content Deviation Metrics
The Content Deviation Metrics assess the deviation of the news content across multiple dimensions by quantifying changes in specific aspects of the news. FUSE-EVAL evaluates the news content based on six key dimensions. (1) Sentiment Shift (SS) measures the change in emotional tone between the original news content and its evolved version (Lu et al., 2022; Ma et al., 2021). Sentiment plays a crucial role in how information is perceived and shared; shifts in sentiment can indicate the introduction of bias or emotional manipulation. (2) New Information Introduced (NII) assesses the extent to which additional information, not present in the original news, has been incorporated during the evolution process (Arif et al., 2016; Wang et al., 2017). The introduction of new facts, details, or claims can significantly alter the original message, potentially leading to misinformation if the new information is inaccurate or misleading. (3) Certainty Shift (CS) evaluates changes in the level of confidence or assertiveness expressed in the news content (Krafft and Spiro, 2019; Kim and Yoon, 2022). Shifts from definitive statements to speculative language (or vice versa) can influence the perceived credibility and urgency of the information. (4) Stylistic Shift (STS) examines changes in writing style, tone, and linguistic features (Wu et al., 2024; Alsaif and Aldossari, 2023). Alterations in style—such as moving from formal to informal language, changing sentence complexity, or using different rhetorical devices—can affect readability and audience engagement. (5) Temporal Shift (TS) measures changes related to time references within the news content (Shen et al., 2024; Mu et al., 2023). Modifications to dates, times, or the sequence of events can significantly impact the interpretation of the news. (6) Paraphrasing Degree (PD) evaluates the extent to which the content has been rephrased or paraphrased from the original text, which may obscure the original meaning or introduce ambiguity.
We employ LLMs to automate the evaluation process of the FUSE-EVAL metrics. Each of these dimensions is scored by GPT-4o-mini on a scale from 1 to 10, where a score of 1 indicates minimal or no deviation from the original news, and a score of 10 signifies a significant or complete deviation along that dimension. By assigning numerical scores, we can quantitatively compare the extent of changes across different aspects of the news content. As shown in the Figure 3, FUSE-EVAL demonstrates cumulative deviations (Pröllochs and Feuerriegel, 2023) during fake news evolution, confirming its effectiveness. To evaluate the overall deviation, we calculate the Total Deviation (TD) for each agent at each time step by averaging the scores:
(5) $\text{TD}_{i}^{t}=\frac{1}{6}\sum_{d=1}^{6}D_{i,d}^{t},$
where $D_{i,d}^{t}$ is the score of dimension $d$ for agent $i$ at time $t$. The detailed evaluation process are provided in Appendix B.
4.2. Statistical Deviation Metrics
The Statistical Deviation Metrics are derived from the Total Deviation scores and provide insights into the overall trends and patterns of news evolution within the network. We then analyze several statistical metrics based on the TD. The $\Delta$Deviation represents the difference in Average Deviation between the final and initial day of the simulation, highlighting the overall growth in deviation throughout the simulation period. The Average Deviation is the mean of TD across all agents at each time step, providing insight into the general trend of news evolution within the network. The Deviation Variance measures the statistical variance of TD among agents, indicating the degree of dispersion and how uniformly the news content deviates across the network. The Final Deviation is the average TD at the final simulation time step $t$, representing the cumulative effect of news evolution over the entire simulation duration. The Maximum Deviation and Minimum Deviation refer to the highest and lowest average TD observed during the simulation, respectively. These metrics indicate the points at which the news content has deviated most and least significantly from the original. The Peak Deviation Time indicates the percentage of the total simulation time taken to reach the Peak Deviation Rate. It provides insight into how quickly the maximum deviation occurs relative to the entire simulation duration. The Half Deviation Time is the time step $t_{0.5}$ at which the average TD reaches half of the Max Deviation, indicating the speed at which significant deviation occurs within the network.
4.3. Implementation Details
Our simulation framework was developed using Python scripts, leveraging various libraries to model the agents and their environment effectively. The LLM used is gpt-4o-mini, accessed via OpenAI API calls. When creating the network structure, we used the Python library networkx to construct different social network structures. The simulation includes 40 agents, whose traits were based on the Big Five personality dimensions commonly used in psychology (Barrick and Mount, 1991). Each agent was assigned scores on these traits to introduce variability in behaviors and interactions within the simulation. For further details, please refer to our code. We ensured that our framework is adaptable to different language models and have provided details on API costs and compatibility with other models in Appendix E and F.
Table 1. Comparative analysis of fake news evolution across different settings, including variations in topics, social networks, spread traits, and intervention strategies. Upward or downward arrows represent better control of fake news evolution.
Comparison Factors Setting $\Delta$ Deviation $\downarrow$
Average
Deviation $\downarrow$
Deviation
Variance $\downarrow$
Max
Deviation $\downarrow$
Min Deviation
Final
Deviation $\downarrow$
Peak Deviation
Time $\uparrow$
Half $\Delta$ Deviation
Time $\uparrow$
Topic Politics 3.148 6.594 0.511 7.440 3.442 6.590 0.133 0.033
Science 1.446 3.533 0.207 4.236 2.026 3.472 0.767 0.033
Network Structure Random 1.905 3.315 0.347 4.206 1.892 4.206 1.000 0.233
Scale-Free 2.631 4.287 0.725 5.652 1.492 4.955 0.767 0.167
High-Clustering 4.313 6.193 1.027 7.030 2.348 6.661 0.500 0.033
Spread Type Normal Spread 1.176 3.536 0.606 4.705 1.398 3.524 0.800 0.133
Emotional Spread 1.688 4.182 0.456 5.105 2.008 4.303 0.333 0.067
Super Spread 2.920 4.434 0.672 5.613 2.054 5.067 0.700 0.100
Traits Impressionable 3.088 4.998 0.956 6.428 2.262 5.677 0.667 0.133
Vigilant 1.945 4.081 0.446 5.021 2.485 4.593 0.400 0.133
Intervention No Intervention 3.208 5.546 1.247 7.340 1.841 6.383 0.767 0.167
Intervention 1.384 4.207 0.476 5.302 1.841 4.559 0.200 0.067
Figure 3. (a) The FUSE-EVAL scores show cumulative deviations of the information over time. (b) Ablation study showing the effectiveness of hierarchical memory and propagation roles. (c) Impact of removing different agent types on fake news evolution. (d) Effectiveness of early-intervention, showing a clear reduction in deviation over time compared to the no-intervention condition.
5. Validation of the FUSE Framework
In this section, we demonstrate the effectiveness of our FUSE framework by showing that FUSE aligns with the influence of certain factors on fake news propagation, reproduces these patterns, and can also replicate real-world fake news.
5.1. Alignment with Real-World Patterns
5.1.1. Topic Comparison
We compared the evolution of fake news across five topics: politics, science, finance, terrorism, and urban legends. Table 1 shows that political fake news spreads the fastest, with average deviation peaking rapidly within four days and remaining high. Terrorism follows a similar pattern. In contrast, financial and science-related fake news spread more slowly, with science-related misinformation evolving the slowest and average deviation remaining consistently low. Table 1 shows the final deviation for political news is approximately 90% higher than that of science news. These results indicate that political fake news is more prone to rapid distortion and widespread belief, while science-related misinformation spreads more cautiously, consistent with prior research (Lazer et al., 2018b).
5.1.2. Social Network Comparison
Using a terrorism topic, we compared fake news evolution across different social network structures: random, scale-free, and high-clustering networks. Table 1 shows that high-clustering networks exhibit the fastest and most extensive spread of misinformation, with deviation peaking rapidly and remaining high. This indicates that tightly connected communities are particularly susceptible to rapid belief distortion, aligning with the ”echo chamber” effect (Cinelli et al., 2021). Random networks show the slowest evolution of fake news, with lower variance suggesting individuals are less likely to form polarized beliefs. Scale-free networks are intermediate. Peak deviation time is the longest in random networks and shortest in high-clustering networks, illustrating that clustering accelerates fake news evolution, consistent with prior research (Lind et al., 2007; Trpevski et al., 2010).
5.1.3. Spread Type Comparison
Using a terrorism topic, we compared fake news evolution across three spread types: normal spread, emotional spread, and super spread. Normal spread follows standard settings, super spread was assigned to high-degree nodes, and emotional spread involves participants using heightened emotional language. Table 1 shows that super spread leads to the highest level of misinformation, due to highly connected influencers amplifying dissemination. The emotional spread has a moderate effect, while normal spread shows the slowest evolution. Super spread reaches peak deviation fastest, followed by emotional spread; normal spread takes the longest. These results indicate that super spread and emotional spread drive faster and more significant misinformation evolution.
5.1.4. Personality Traits Comparison
Using a terrorism topic, we compared the impact of personality traits on fake news evolution. Based on the Big Five personality traits (Barrick and Mount, 1991), individuals with high agreeableness and neuroticism (Impressionable) are more likely to believe rumors, while those with low levels (Vigilant) are less susceptible. Table 1 show that impressionable agents are more prone to accepting and spreading misinformation. Vigilant agents exhibit stronger resistance, with beliefs remaining relatively stable. These results align with previous studies and demonstrate our framework’s effectiveness in simulating the influence of personality traits on fake news spread.
5.2. Alignment with Real-World Fake News
We conducted experiments across various topics and found that the fake news evolved by the FUSE framework closely corresponds to real-world fake news. As shown in Figure 4, the news about ”Trump being attacked” starts as true, evolves into partially false, and eventually becomes entirely fake. The agent, as a commentator, often adds its own views, while its neighboring verifiers and spreaders act according to their roles. Additionally, our framework generates fake news such as ”Trump was not attacked, it’s a dramatic effect,” which is also a widely circulated piece of fake news in the real world [^2] https://x.com/cwebbonline/status/1814708054916784594333https://x.com/EndWokeness/status/1813898763100176484. From a quantitative analysis perspective, for each topic, 73% of fake news is recovered by our framework. The detailed case study and analysis results are provided in the Appendix C.
6. ANALYSIS AND DISCUSSION
6.1. Ablation Study
We chose a terrorism topic to demonstrate the effectiveness of our model’s components and conducted two ablation studies to evaluate the contribution of key components in the FUSE framework.
6.1.1. The Impact of Hierarchical Memory and Propagation-Role.
As shown in Figure 3(b), the full FUSE framework demonstrates a clear accumulation effect in deviation, indicating its effectiveness in simulating fake news evolution. After removing hierarchical memory, the deviation significantly drops, with a 39.8% reduction throughout the simulation. The absence of the accumulation effect indicates the simulation fails (Pröllochs and Feuerriegel, 2023). This highlights the critical role hierarchical memory plays in capturing the depth and persistence of belief distortion, as it enables agents to store and reflect on both short-term and long-term information. Similarly, removing the propagation role further decreases the deviation. This underscores the importance of agent roles (such as spreader, commentator, verifier, and bystander) in shaping the propagation and evolution of information. Without these roles, the agents behave more uniformly, and the accumulation effect of deviation disappears, meaning that the news does not evolve.
6.1.2. The Impact of Propagation Role Types.
After the first ablation study showed that removing the propagation role leads to simulation failure, we conducted a second study to assess the impact of different role types (spreader, commentator, verifier, and bystander) on fake news evolution. As shown in Figure 3(c), removing the commentator agent had the most significant impact, with a sharp drop in average deviation, leading to simulation failure. This confirms the central role of the commentator in amplifying and disseminating misinformation. In contrast, removing the spreader slightly reduced the deviation, as spreaders do not provide explanations or opinions, which are key drivers of news distortion but may still contribute to its evolution. Interestingly, removing the verifier increased deviation, as verifiers play a crucial role in checking the accuracy of information. Without verifiers, the system becomes more susceptible to misinformation, leading to a greater spread of fake news. Finally, removing the bystander had the least impact, as bystanders passively observe without actively participating in the information spread.
Overall, these ablation studies demonstrate the effectiveness of FUSE. Hierarchical memory, propagation roles, and specific agent types play important roles in shaping the evolution of fake news within the FUSE framework.
6.2. Fake News Intervention
Figure 4. A Case of FUSE: A true news gradually evolves into partially false and eventually entirely fake news over time.
6.2.1. Intervention Strategy.
Based on the previous experimental results, we chose to introduce the official agent at high-degree nodes. Table 1 presents the comparative experimental results between the no-intervention and intervention conditions. Since fake news evolution reached a high point on the sixth day, we implemented the intervention strategy, where the official agent released a statement. The deviation under the intervention condition decreased by 37.8% compared to the no-intervention condition. This early intervention effectively blocked the initial spread of fake news, preventing a large number of agents from believing in the misinformation. However, as the news continued to propagate, the effectiveness of the intervention gradually weakened. By the 12th day, although the deviation under the intervention condition was still lower than that under the no-intervention condition, the gap narrowed to 22.3%, suggesting that the impact of a single intervention diminishes over time, as agents continue to interact and may revert to their previous beliefs or encounter conflicting information. Around the 16th day, we conducted a second intervention, which again reduced the deviation by 31.8% compared to the no-intervention condition. This further demonstrates the importance of continuous and timely fact-checks to suppress the spread of fake news. Compared to the no-intervention condition, the final deviation under the intervention strategy was reduced by approximately 28.6%  and the variance in deviation was also reduced by 61.8%. Moreover, the peak deviation time under the intervention strategy occurred approximately 0.56-time units earlier than in the no-intervention case.
Overall, the intervention strategy consistently maintained a lower average deviation than the no-intervention condition. The results highlight the importance of not only introducing early interventions but also maintaining regular interventions over a longer period to effectively combat the continuous spread and evolution of fake news.
6.3. Factors in Fake News Evolution
Figure 5. (a) The contribution percentages of factors in FUSE-EVAL to fake news evolution. (b) Comparison of the contributions of these factors across different topics, with Politics and Terrorism showing balanced contributions, while Science relies more on NII and less on TS and STS.
The analysis of experimental results and charts indicates varying contributions of different factors to fake news evolution. The Figure 5 (a) shows that PS contributes the most (22.3%), suggesting that altering reporting angles or distorting original information is the key driver of fake news evolution. NII follows with 18%, highlighting its significant role in this process. SS an STS contribute 17.2% and 17.5%, respectively, while TS has the smallest impact at 11%. The Figure 5 (b) reveals topic-specific patterns. Political and terrorism-related fake news evolves across multiple dimensions, especially new information, perspective, and sentiment shifts. In contrast, science-related fake news is driven mainly by new information, with less influence from temporal or style shifts. Urban legends and finance topics also heavily rely on perspective shifts and new information. In summary, PS and NII are the main drivers of fake news evolution, with time-related changes having the least impact. Understanding these patterns can help in developing targeted strategies to detect and mitigate misinformation.
7. CONCLUSION
We introduced the FUSE framework, a simulation environment that models the evolution of true news into fake news using agent-based modeling and LLMs. By implementing the FUSE-EVAL evaluation framework, which quantifies content deviation across six dimensions, we systematically analyzed how misinformation propagates and transforms within social networks. Our experiments demonstrated alignment with real-world observations across various settings, reinforcing existing theories on misinformation spread, such as the rapid distortion of political news, the influence of network clustering, the impact of super spreaders and emotional content, and the role of individual personality traits in susceptibility to fake news. By automating the evaluation process with LLMs like GPT-4o-mini, our framework achieves scalable and consistent analysis. By enhancing our understanding of misinformation spread, our framework supports the development of a more responsible and trustworthy web environment.
Appendix A Prompt Set
Here, we present a detailed description of the prompts employed in our FUSE framework to model the dynamics of fake news evolution.
1. The prompt for the role-specific reintroduction function $f_{r_{i}}$ is as:
$f_{spr}$: share information quickly without verifying its accuracy. $f_{com}$: modifies or adds their views before sharing news. $f_{ver}$: performs some verification before spreading news. $f_{bys}$: consume news without participating in its dissemination.
2. The prompt for Short-Term Memory function $f_{S}$ is as:
Summarize the opinions you have heard in a few sentences, including their own perspective on the news.
3. The prompt for Long-term memory function $f_{L}$ is as:
Review the previous long-term memory and today’s short-term summary. Please update the long-term memory by integrating today’s summary, ensuring continuity and incorporating any new insights.
4. The prompt for the reasoning function is as:
As a [role], you combine your [previous personal opinion] with the new information stored in your [long memory]. You process this information in the following manner: [role behavior], and then reintroduce the [news].
5.The prompt for ”Official Statement” is as:
According to the current investigation, That [news] is true. We have noticed that some social media platforms and certain media outlets are spreading false information, claiming that [news] is fake. We firmly state that such claims are baseless. The government is committed to transparency and will provide timely updates on the investigation. We urge the public to seek accurate information from official channels, and necessary actions will be taken against those who intentionally spread false information.
Appendix B Human Evaluation
To efficiently evaluate the deviation of news content across the multiple dimensions defined in FUSE-EVAL, we employ large language models (LLMs) to automate the assessment process. This approach provides consistent and scalable evaluations, reducing the reliance on time-consuming human evaluation. We utilize two versions of OpenAI’s language models: gpt-3.5-turbo and GPT-4. For each agent’s news content at various time steps, we prompt the LLMs to evaluate the six FUSE-EVAL dimensions by comparing the evolved content with the original news article,which is as follows:
  • Sentiment Shift (SS)
  • New Information Introduced (NII)
  • Certainty Shift (CS)
  • Stylistic Shift (STS)
  • Temporal Shift (TS)
  • Perspective Deviation (PD)
Table 2. Correlation for LLM-based evaluations across FUSE-EVAL dimensions.
Dimension GPT-3.5-turbo GPT-4o-mini
Sentiment Shift (SS) 0.524 0.705
New Information Introduced (NII) 0.621 0.765
Certainty Shift (CS) 0.527 0.719
Stylistic Shift (STS) 0.481 0.642
Temporal Shift (TS) 0.503 0.694
Perspective Deviation (PD) 0.548 0.760
Average Correlation 0.531 0.714
The prompt used is as follows:
I have an original news and multiple related news. I want to evaluate how much these news deviate from the original news based on the following criteria: 1. Sentiment Shift: How does the sentiment of the news compare to the original news? Is the tone more positive, negative, or neutral compared to the original? 2. Introduction of New Information: Does the news introduce additional information not in the original news, such as political conspiracy or speculation? Evaluate how much of the article is focused on these new details. 3. Certainty Shift: How does the news language change in terms of certainty? Does it use more ambiguous terms like “possibly” or “allegedly” compared to the original news, or does it present the information with more certainty? 4. Stylistic Shift: How does the writing style compare to the original? Has the news moved from neutral reporting to a more exaggerated or dramatic tone? 5. Temporal Shift: Does the news shift focus from the specific event mentioned in the original news to broader or unrelated timeframes, such as mentioning legal battles or long-term political issues? 6. Perspective Deviation: Does the article introduce subjective opinions or perspectives that deviate from the objective reporting in the original news? For instance, questioning the truth of the event or speculating on hidden motives. Task: Please evaluate the following news based on each criterion and provide a score from 0 to 10, where 0 means the article is completely aligned with the original news, and 10 means it has fully deviated. Original News:[original news] News articles to Evaluate:[Evolved News] Please provide the results in the following format: [output format]
The models assign scores from 1 to 10 for each dimension based on predefined evaluation criteria.
To validate the effectiveness of using LLMs for this task, we conduct a benchmarking study by comparing the LLM-generated evaluations with those from human judges. A group of human evaluators, knowledgeable in journalism and communication studies, independently assess a representative sample of the news content using the same scoring guidelines. We calculate the Pearson correlation coefficients between the scores assigned by the LLMs and the human evaluators for each dimension. The results, summarized in Table 2 , demonstrate that GPT-4o-mini exhibits a high correlation with human evaluations across all dimensions, outperforming gpt-3.5-turbo.
The high correlation coefficients indicate that GPT-4o-mini closely aligns with human evaluations, making it a reliable tool for assessing news deviation in our simulation. We achieve a scalable and consistent assessment process by leveraging GPT-4o-mini for evaluation. This approach allows us to efficiently analyze large volumes of data generated in the simulation while maintaining evaluation quality comparable to human judgments. The strong alignment with human evaluations validates using GPT-4o-mini as an effective evaluator of news content deviation across the FUSE-EVAL dimensions.
Appendix C Alignment Between Simulated and Real-World Fake News
Additionally, our framework generates fake news narratives that closely mirror those found in the real world. This alignment validates the realism of our simulation and demonstrates its potential as a tool for studying misinformation dynamics. By producing content that reflects actual fake news, our framework enables researchers to better understand how such information originates and spreads, thereby aiding in the development of effective strategies to combat misinformation.
The specific case is as follows:
Appendix D Various Topics and Simulation Results
Figure 6. The average deviation of news changes across different topics, social networks, dissemination role types, and traits.
In our experiments, we compared the evolution of fake news across five different topics: politics, science, finance, terrorism, and urban legends. As shown in Figure 6(a), political fake news spreads the fastest, with average deviation rapidly peaking within just four days and remaining at a high level. Fake news related to terrorism follows closely behind, showing similarly fast spread, likely due to the emotional intensity and urgency associated with such topics, which prompt individuals to quickly form beliefs and propagate the news widely. In contrast, financial news spreads at a slower pace, with deviation gradually accumulating over time. Although financial news is significant in terms of economic impact, individuals tend to engage in more rational thinking when encountering such news, leading to more stable growth in average deviation. Science-related fake news evolves the slowest, with average deviation consistently remaining low throughout the propagation process. These results is consistent with previous studies (Lazer et al., 2018a). This suggests that individuals are generally more cautious when dealing with scientific topics, often subjecting the information to more thorough verification.
Here, we provide detailed descriptions of the news items used in our experiments on fake news evolution across various topics.
Political - In 2023, the Argentine government announced a new debt restructuring agreement with the International Monetary Fund (IMF), accepting a series of austerity measures in exchange for a new round of loan assistance. Science - The discovery and successful use of CRISPR-Cas9 technology to edit genes in animals. Scientists have made breakthroughs in curing genetic disorders in mice, opening doors for future human treatments. Terrorism - Trump was attacked at a campaign rally in Butler, Pennsylvania, and eyewitnesses say his ear was injured. Urban Legends - Stella Liebeck was awarded damages after suffering third-degree burns from spilled coffee. Finance - The Bernie Madoff Ponzi scheme, which collapsed in 2008, defrauded investors of billions of dollars.
Additionally, we demonstrate the effectiveness of our FUSE framework by showing that it aligns with the influence of various factors, including social network structure, type of propagation, and agent traits, on the evolution of fake news. FUSE reproduces these patterns and can also replicate real-world fake news dynamics, as illustrated in Figure 6.
Appendix E Analysis of Experimental Costs
In this section, we analyze the costs associated with our experiments utilizing the GPT-4o-mini APIs. At the time of our experiments, OpenAI’s pricing model was as follows: for gpt-4o-mini, the cost was 0.15 USD for every 1M input tokens and 0.6 USD for every 1M output tokens.
Our simulations involved multiple agents interacting over several days, with each agent generating and processing textual content. For a simulation with 40 agents over 30 days, it involved approximately 3 to 5M input tokens and 5 to 10M output tokens. This resulted in an estimated cost of 4 USD to 8 USD for the entire simulation phase using gpt-4o-mini combining both the simulation and evaluation phases.
Conducting comparable research in real-world settings typically involves significantly higher expenses. Real-world studies require funding for participant recruitment, compensation, data collection tools, infrastructure setup, and extended durations to gather and analyze data. Depending on the scale and scope, such studies can cost from several thousand to hundreds of thousands of dollars. By leveraging GPT-4o-mini, we can simulate complex social interactions and the evolution of information without the logistical challenges and high costs associated with real-world experiments. This approach allows for rapid iteration and scalability, enabling us to explore various scenarios and intervention strategies efficiently. This cost analysis highlights the economic advantages of our simulation-based methodology-FUSE. The ability to conduct extensive experiments at a fraction of the cost demonstrates the practicality and accessibility of using LLMs for research in misinformation dynamics. It opens avenues for researchers with limited resources to contribute valuable insights into the field, fostering a more inclusive and innovative research environment.
Social networks in real life can generally be categorized into three types: high clustering networks, scale-free networks, and random networks, which correspond respectively to Figure 7 (a), (b), and (c).
Appendix F Simulation on Different Backbones
Figure 7. Different social network in our framework: (a) high clustering network (b) scale-free network (c) random network
Figure 8. Average Deviation changes with GPT-4 and GPT-4o-mini as the backbone under the terrorism topic, both of which demonstrate a deviation accumulation effect.
To further validate the robustness and adaptability of our FUSE framework, we conducted additional experiments using different LLMs as the backbone. Specifically, we implemented simulations with both GPT-4o-mini and GPT-4 to assess whether the choice of LLM affects the effectiveness of our framework.
As shown in Figure 8, in simulations focused on political topics, we observed that when using GPT-4 as the underlying LLM, the number of agents adopting and spreading misinformation increased rapidly. This surge led to a majority of agents holding and propagating distorted versions of the original news. Notably, this pattern was consistent with the results obtained when GPT-4o-mini was used as the backbone, indicating that the dynamics of misinformation spread are preserved across different LLMs. These consistent results demonstrate that our FUSE framework effectively captures the core mechanisms of fake news evolution and public opinion formation, independent of the specific LLM used to power the agents.
By showing that FUSE performs effectively with different LLM backbones, we confirm that the framework is not only robust but also adaptable to various technological settings. This adaptability is particularly valuable given the rapid development of LLM technologies, ensuring that our framework remains relevant and effective as newer models become available. In summary, the consistent performance of our simulation across different LLMs underscores the effectiveness of the FUSE framework in modeling misinformation propagation. It highlights the framework’s potential for broad application in studying fake news dynamics and developing strategies for mitigation, regardless of the underlying language model technology.
Appendix G Social Network
High clustering networks are characterized by nodes that tend to form tightly knit groups or communities, where neighbors of a node are likely to be neighbors themselves. The degree of clustering can be quantified by the clustering coefficient $C$, which is defined for a node $v$ as:
$C_{v}=\frac{2T(v)}{k_{v}(k_{v}-1)},$
where $T(v)$ is the number of triangles passing through node $v$ and $k_{v}$ is the degree of $v$. The clustering coefficient for the whole network is the average of $C_{v}$ over all nodes $v$.
Scale-free networks are characterized by a power-law degree distribution, where the probability $P(k)$ that a randomly selected node has $k$ connections to other nodes follows:
$P(k)\sim k^{-\gamma},$
where $\gamma$ is a parameter typically in the range 2 ¡ $\gamma$ 3. This distribution implies that most nodes have few connections, while a few hub nodes have a large number of connections. This heterogeneity in node connectivity is a hallmark of scale-free networks.
Random networks, also known as Erdős–Rényi networks, each edge is included in the network with a fixed probability $p$ independent of the other edges. For a network with $n$ nodes, the probability $P(k)$ that a randomly selected node has $k$ connections is given by the binomial distribution:
$P(k)=\binom{n-1}{k}p^{k}(1-p)^{n-1-k}.$
For large $n$, this can be approximated by the Poisson distribution:
$P(k)\approx\frac{\lambda^{k}e^{-\lambda}}{k!}$
where $\lambda=p(n-1)$ is the expected degree of a node. These three types of networks are used in the environment simulation of news evolution within our FUSE framework.